-15x+10-3x=-6x+8-4x
plz solve for x!

Answer by Boreal(15235)

-15x+10-3x=-6x+8-4x
Collect terms
-18x+10=-10x+8
add 18x to both sides
10=8x+8
subtract 8 from both sides
2=8x
divide by 8 both sides
(1/4)=x
check
3.75+10-0.75=-1.5+8-1
5.5=5.5